OpenStudy (anonymous):

ALWAYS FAN AND MEDAL A lawn sprinkler rotates 60 degrees to water the grass. If the sprinkler reaches 15 feet away , what is the area of the sector of grass that gets watered?

OpenStudy (anonymous):



OpenStudy (anonymous):

total area of the circle is \(\pi r^2=\pi\times15^2\)

OpenStudy (anonymous):

what portion do you have?

OpenStudy (anonymous):

if the answer is not clear, recall that there are a total of 360 degrees in the circle

OpenStudy (anonymous):

everything that's up there. Is 15 the radius?

OpenStudy (anonymous):

yes 15 is the radius

OpenStudy (anonymous):

\[15^{2}= 225\]

OpenStudy (anonymous):

ok but you don't have the full circle whose area is \(225\pi\) you only have a portion of it

OpenStudy (anonymous):



OpenStudy (anonymous):

\[225\times \pi = 706.8583470577\]

OpenStudy (anonymous):

but, i repeat, that is the entire circle , you only have a part of it

OpenStudy (anonymous):

But its only asking for the area of that specific portion of grass